OpenStudy (anonymous):

Quadrilateral ABCD is inscribed in a circle. m angle A is 64°, m angle B is (6x + 4)°, and m angle C is (9x - 1)°. What is m angle D?

OpenStudy (anonymous):

82

OpenStudy (raden):

A + C = 180 degrees 64 + 9x - 1 = 180 solve for x, first

OpenStudy (anonymous):

180-82 = 98 answer: 98

OpenStudy (raden):

yes, correct!

OpenStudy (anonymous):

thank you!!